OpenStudy (anonymous):

How to you find the angle measure of m<9 and m<10? m<9=2x-4 m<10=2x+4

jimthompson5910 (jim_thompson5910):

How are angles 9 and 10 related? Are they supplementary? Complementary? If there's a picture, please post it

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

ok so they are supplementary

jimthompson5910 (jim_thompson5910):

this means that m < 9 + m < 10 = 180 (2x-4) + (2x+4) = 180

jimthompson5910 (jim_thompson5910):

Tell me what you get when you solve for x

OpenStudy (anonymous):

45

jimthompson5910 (jim_thompson5910):

good, now plug this into 2x-4 to get what?

OpenStudy (anonymous):

86

jimthompson5910 (jim_thompson5910):

so that's the measure of angle 9

jimthompson5910 (jim_thompson5910):

do the same for 2x+4 to find the measure of angle 10

OpenStudy (anonymous):

94

jimthompson5910 (jim_thompson5910):

good, so angle 9 is 86 degrees and angle 10 is 94 degrees

OpenStudy (anonymous):

Thank you : )

jimthompson5910 (jim_thompson5910):

you're welcome